OpenStudy (anonymous):

lim x-> infinity e^x / ln(1-2/x)

OpenStudy (anonymous):

Could try L'Hopital's rule here.

OpenStudy (anonymous):

i am suppose to do L'hospital rule..

OpenStudy (anonymous):

Are you having trouble finding the derivatives then?

OpenStudy (anonymous):

and then i dont even know what i was doing..i can show u my work if u want..but i dont think u can understand it ..

OpenStudy (anonymous):

Let's just go one step at a time: L'Hopital's rule says to differentiate the top and bottom of the fraction. Find the derivative of \(e^x\) and the derivative of \(ln(1-2/x)\).

OpenStudy (anonymous):

yea, dont u need to first prove it is 0/0 or infinity / infinity

OpenStudy (anonymous):

Not necessarily.

OpenStudy (anonymous):

L'Hopital's rule will work regardless; it's just that it is most useful in those indeterminate cases.

OpenStudy (anonymous):

Right now it is ∞/0, which is just as bad.

OpenStudy (anonymous):

yea, i can still apply l'hopital rule even it is infinity/ 0?

OpenStudy (anonymous):

As far as I know, L'Hopital's rule always works regardless.

OpenStudy (anonymous):

derivative is (1/ 2-(2/x) ) (-2/x^2) / -e^2x

OpenStudy (anonymous):

e^-x that's becuz i bring e^x down to get 0/0

OpenStudy (anonymous):

it's #16

OpenStudy (anonymous):

That is not making sense to me . . . This is what you are trying to find, right? \(\large \lim_{x \rightarrow ∞} \frac{ e^x}{ ln(1-2/x)} \)

OpenStudy (anonymous):

yea..

OpenStudy (anonymous):

The derivative of \(e^x\) is \(e^x\) and the derivative of \(ln(1-2/x)\) is \(\large \frac{2x^{-2}}{1-2x^{-1}}\)

OpenStudy (anonymous):

Does that check out to you?

OpenStudy (anonymous):

yea, that is what i got too, it just looks messy ..

OpenStudy (anonymous):

It can be cleaned up a little by rearranging things. If you put the derivative of the top over the derivative of the bottom, you should be able to see that it tends to -∞

OpenStudy (anonymous):

i actually cant see it is going to -ve infinity..

OpenStudy (anonymous):

2/ infinity is 0 isnt it

OpenStudy (anonymous):

It can be rearranged to \((x^2e^x-2xe^x)/2)\)

OpenStudy (anonymous):

urgg..no idea how u do it

OpenStudy (anonymous):

Let's see.. Derivative of e^x is e^x, derivative of ln(1-2/x) is 2/(x^2-2x). Verify these two things first. One over the other makes (e^x)(x^2-2x)/2. You can then distribute the e^x through the parentheses.

OpenStudy (anonymous):

isnt derivative of ln(1-2/x) = 2x^-2 / 1-2x^-1

OpenStudy (anonymous):

Yes, but you can simplify it using rules for exponents.

OpenStudy (anonymous):

i actually dont know how to do that..would u mind showing me how to do it :/

OpenStudy (anonymous):

x^-2 means that it is x^2 in the denominator. so \(\large \rightarrow \frac{2}{x^2( 1-2x^-1)}\) You can then distribute the x^2 through the parentheses.

OpenStudy (anonymous):

oh ok got it thx